The ball is tied to a string to whirl it having a radius of 40.0 cm at a velocity of 2.5 m/s. What will be



the acceleration of the ball?

1
Expert's answer
2022-03-26T15:31:39-0400

The acceleration of the ball will be


"a=v^2/r=2.5^2/0.4=15.6\text{ m/s}^2."